40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Dec 27 05:38:56 2021, rana, Update, Control System, cavity locking loop shape 
    Reply  Mon Jan 24 10:24:49 2022, shruti, Update, Control System, cavity locking loop shape OLTF_fit.pdf
       Reply  Wed Jan 26 17:28:53 2022, shruti, Update, Control System, cavity locking loop shape ctrl_pdh_5kHz_Screenshot.pngctrl_pdh_10kHz_Screenshot.pngscope_red.pdffilter_red.pngctrl_pdh_100kHz_Screenshot.png
          Reply  Mon Jan 31 19:03:43 2022, shruti, Update, Control System, cavity locking loop shape - pomona box TF FilterTF.pdfdata_and_analysis.zip
Message ID: 2843     Entry time: Wed Jan 26 17:28:53 2022     In reply to: 2841     Reply to this: 2849
Author: shruti 
Type: Update 
Category: Control System 
Subject: cavity locking loop shape 

Some observations while measuring the noise spectra [Rana, Shruti]:

  • The cantilever mode at 45 Hz is the dominant source of noise in the measured noise spectra. Attachment 1 shows the noise spectrum at lower frequencies and Attachment 3 shows this signal in the PDH signal, and cavity transmission and reflection.
  • The contribution of the noise of the second order (90 Hz) is roughly 35 dBV smaller than the fundamental, and the third harmonic at 135 Hz seems virtually non-existent visually.
  • There are present many peaks from 1 kHz  to tens of kHz. The smaller spacing between them seems to be around the cantilever resonance. (Attachment 2). Although very similar looking peaks do seem to be present in noise spectra measured before we added the cantilever (such as in elog 2776).

In order to suppress these noisy peaks, following the suggestion in elog 2838, I made the pomona box filter shown in attachment 4. The R_{in} is the output impedance of the LB1005 servo and R_{out} = 10 k\Omega is the input impedance of the analog modulation port of the ITC502 current driver being used.

[Update 31st Jan 22] R_{in} is shown incorrectly in Attachment 4 - it is supposed to be in series, not in parallel to the remaining circuit.

Attachment 1: ctrl_pdh_5kHz_Screenshot.png  878 kB  Uploaded Thu Jan 27 14:02:56 2022  | Hide | Hide all
ctrl_pdh_5kHz_Screenshot.png
Attachment 2: ctrl_pdh_10kHz_Screenshot.png  1.258 MB  Uploaded Thu Jan 27 14:03:06 2022  | Hide | Hide all
ctrl_pdh_10kHz_Screenshot.png
Attachment 3: scope_red.pdf  146 kB  Uploaded Thu Jan 27 14:20:21 2022  | Hide | Hide all
scope_red.pdf
Attachment 4: filter_red.png  418 kB  Uploaded Thu Jan 27 14:22:49 2022  | Hide | Hide all
filter_red.png
Attachment 5: ctrl_pdh_100kHz_Screenshot.png  1.372 MB  Uploaded Thu Jan 27 14:31:11 2022  | Hide | Hide all
ctrl_pdh_100kHz_Screenshot.png
ELOG V3.1.3-